InfocomnGameoInformation PaulD.Smith paulsmith@dg.com ThemtablelbelowlcontainslallkInfocommgamesnandkallkknownlinformation abouteachgameandeachrevision.AlltheinformationbelowwascontributedbyPaulDavid Doherty<d.doherty@bamp.berlinet.in-berlin.de>,exceptmostoftheproductcodesanddifficulty ratingsandalltheinformationonLTOIIIfortheIBM-PC,whichwerecontributedbyDaniel Sharpe<dpsharpe@eos.ncsu.edu>,andtheinformationonLTOIfifi&IIrthecintosh,ich waslcontributedlbykmem:-). Themcolumnsninktheltablelandktheirlmeaningsnarelaslfollows: GameTitle Specifiesotheltitlelofkthelgame.Gamesoarellistedminkalphabeticalmorder. Aut (Author)Liststheinitialsoftheauthor(s)ofthegame.Authorsarelistedbelow,along withlwhatltheylweremknownltokbeldoingkinkMarch,k1992: DADouglaslAdamsm-lWorkingkonkthel5thkbookkinkthelHHGttGmseries DAnDavelAndersonm-l? BBBoblBatesn-lPresidentkofkLegendmEntertainment MBeMichaellBerlynm-lWritesmSEGAlgamesnforkAccolade MBMarclBlankl-lOwnsmakComputerLandnstoreminkBend ABAmykBriggsm-lAkPhDkcandidateminkExperimentalkPsychology JCJamesnClavelll-l? SGStukGalleym-lWorkslatkThinkinglMachinesmCorp. JLJimlLawrenceo-l? DLDavidkLeblingl-lWritesmakspreadsheet SMStevelMeretzkyn-lWritesmadventuresmforkLegendmEntertainment BMBrianlMoriartyk-lWorkslatkLucasfilmnGames JONJeffmO’Neilll-l? JWJerrylWolperl-l? DL (Difficulty